# SF-Loc: A Visual Mapping and Geo-Localization System based on Sparse Visual Structure Frames

### v2

@@ VI-B Localization Performance
As to user-side localization, we check the coarse-to-fine localization performance of the proposed system.
@@ VI-B1 Coarse Localization
[…]
Taking the $\Xi=0.4$ case as example, several typical query cases are shown in Fig. 10, which indicate the challenge of re-localization, with dynamic objects, light condition variation and high-ambiguity environmental patterns. In such scenarios, the traditional handcrafted place recognition method (ORB + DBoW2) becomes almost useless, as the low-level appearance of the environment changes significantly for the cross-season sequences. For learning-based methods, the classic NetVLAD achieves 59% 10 m recall, while more modern methods (e.g., CosPlace, EigenPlaces and MixVPR) can generally achieve 90% 10 m recall. Although 90% recall seems good enough for place recognition, it can’t serve as a reliable information source for the localization task, especially for unmanned systems. By naively combining the VPR results of multiple frames, the 10/20 m recall can increase by around 1%. In contrast, with the proposed SSS-based multi-frame VPR technique, the 10/20 m recall can increase by 5%. Especially, for the SSS10 scheme, the 20 m recall can achieve 100%, both for ResNet50 and ResNet18 backbones. Compared to single-frame VPR methods, the introduction of SSS greatly improves the reliability of the coarse localization information, which is crucial for further fine localization.
We also list the results based on a more sparse map with co-visibility threshold $\Xi=0.3$. Generally, the recall is noticeably lower than the $\Xi=0.4$ case, while the EigenPlaces50 (SSS10) scheme can still achieve 93.78%/100% 10m/20m recall.
Figure 11: Horizontal position errors of the fine localization. For SF-Loc, the $\Xi=0.4$ map is used. See TABLE III for detailed description.
@@ VI-B2 Fine Localization
As to the evaluation of fine localization, we test the re-localization modes of ORB-SLAM3 and HLoc as the baselines, and also test GNSS-based schemes for reference, including single point positioning (SPP) and VINS-Fusion (stereo VIO + SPP). For the proposed SF-Loc, we test both single-frame and multi-frame re-localization modes. For multi-frame re-localization, the user-side relative poses are provided by stereo VIO. The factor graph optimization implementation in Sect. V-B is used. For comparison, we also test PnP with RANSAC upon the SF-Loc map.
For ORB-SLAM3 and HLoc, the coarse-to-fine localization is performed based on their own pipeline and map data (corresponding to TABLE I). For SF-Loc, the map data generated with different co-visibility thresholds ($\Xi=0.3/0.4$) are considered, and the coarse localization results of EigenPlaces50 (SSS10) are used.
TABLE III: Accuracy of the Fine Re-Localization. Horizontal Position Errors are Evaluated.
Map Method1 Avail. @0.5 m Avail. @1.0 m Avail. @5.0 m RMSE2 (m)
@@
- VINS-Fusion 0.00% 0.00% 28.98% 6.506
PC (ORB) ORB-SLAM3 0.33% 1.28% 1.55% 0.997
PC (SP) HLoc 88.95% 93.00% 96.39% 0.503
SF (0.4) PnP 70.41% 85.95% 98.17% 0.852
FGO (1) 78.46% 88.78% 98.28% 0.825
FGO (2) 87.23% 94.00% 99.17% 0.604
FGO (5) 96.39% 99.33% 99.89% 0.298
FGO (10) 99.00% 100.00% 100.00% 0.209
FGO (10)* 91.62% 99.67% 100.00% 0.247
SF (0.3) PnP 68.41% 85.01% 96.45% 0.881
FGO (1) 74.68% 86.79% 96.84% 0.807
FGO (2) 86.12% 93.50% 99.22% 0.577
FGO (5) 94.45% 99.22% 100.00% 0.323
FGO (10) 97.89% 100.00% 100.00% 0.213
FGO (10)* 94.73% 99.67% 100.00% 0.248
@@
    2 For RMSE calculation, only $<$ 5 m cases are taken into account.
Figure 12: Examples that demonstrate how multi-frame information contributes to the fine localization. In every panel, the left column corresponds to the SF-Loc (1) scheme, while the right column corresponds to the SF-Loc (2) scheme. In the bird’s eye view plots, the green markers indicate inliers while the red ones indicate outliers.
@@
As can be seen from Fig. 11, ORB-SLAM3 can hardly perform successful re-localization, as both the coarse and fine localization phases are challenged by the significant appearance variation. For HLoc and SF-Loc, decimeter-level localization is achievable when the map matching is well performed, while outliers with several-meter error are also present, which generally correspond to the ambiguous scenarios. It is noted that HLoc performs better than single-frame SF-loc, which is reasonable as the HLoc map contains many more frames (see TABLE I) and stores the fine descriptors extracted from the lossless raw images. In contrast, SF-Loc only uses highly compressed images at the user end. In addition, we witness an improvement of using factor graph optimization than using classic PnP method for the one-frame pose estimation, as the noise model can be better specialized considering the low-resolution depth and the noise of feature matching.
@@
It is impressive that, the localization accuracy of SF-Loc can be greatly improved by integrating multi-frame information. The examples that multi-frame observations lead to better data association and localization results are depicted in Fig. 12. As can be seen, the introduction of multi-frame information not only help bridge the low-correspondence cases, but also help adjust the inlier set of correspondences and improve the geometric configuration, leading to more robust pose estimation. For the $\Xi=0.4$ map, with 10-frame FGO, the SF-Loc achieves 99.0% availability@0.5 m and a RMSE of 0.209 m. As to the absolute positioning accuracy considering the error of map frames, the availability@0.5 m is 91.62% and the RMSE is 0.247 m, which mainly satisfy the demand of decimeter-level localization accuracy. When the map becomes sparser ($\Xi=0.3$), the multi-frame localization performance remains almost unchanged, which shows the possibility of harder map compression in the cost of some redundancy.
@@
Finally, we show the time consumption statistics of user-side localization in TABLE IV, which is tested on a laptop with Nvidia RTX 4080 GPU.
TABLE IV: Time Cost of Different Modules in User-Side Localization.
Phase Module Time cost (ms)
Coarse Localization 1 VPR Descriptor 5.7
Retrieval 27.3
Fine Localization 2 Local Feature Extraction 33.0
Feature Matching 30.3
Pose Estimation 32.2
  • 1 Using “EigenPlaces50 (SSS10)” scheme in TABLE II.
    2 Using “SF-Loc (10)” scheme in TABLE III.
@@ VII Conclusion
In this paper, we propose a vision-centered mapping and localization framework based on the map representation of visual structure frames. The highlight of the system is the lightweight map storage by making the frame data compact and keeping the sparsity, while can still support high-recall, high-accuracy localization through the utilization of user-side multi-frame information.
In later work, we will investigate the application of the system to indoor environments, and improve the system performance by introducing the power of visual foundation models.
Acknowledgments
The implemented SF-Loc is developed by the GREAT Group, School of Geodesy and Geomatics, Wuhan University. The numerical calculations in this paper have been done on the supercomputing system at the Supercomputing Center of Wuhan University.

### v3

@@ VI-B Localization Performance
As to user-side localization, we check the coarse-to-fine localization performance of the proposed system. Considering realistic cases of localization query, we test single-frame and multi-frame (with temporal fusion) localization performance, while historical prior information (like Kalman filtering) is not considered.
@@ VI-B1 Coarse Localization
The test of coarse localization is based on the map frames generated by SF-Loc. The coarse localization performance directly affects further fine-grained association, and a <<20 m recall distance is generally needed for possible fine localization under the experimental scenario. We show the recall rates of different schemes in TABLE II, including the classic ORB + DBoW2 implementation[8] and DNN-based methods (NetVLAD[27], AnyLoc[31], MixVPR[45], CosPlace[29], EigenPlaces[30]). For CosPlace and EigenPlaces, different backbones (ResNet18 or ResNet50[47]) are considered. Based on the state-of-the-art lightweight model EigenPlaces, we employ the proposed SAS technique to aggregate the information of the current and last $K$ sequential queries, termed as “EigenPlaces (SASKK)”, where different window lengths are considered. To verify the effectiveness of the proposed SAS metric, we also implement a naive multi-frame VPR model by clustering the top-$10$ candidates of the current and last $K$ sequential queries and choosing the candidate belonging to the largest cluster as the result, termed as “EigenPlaces (KK)”.
Taking the $\Xi=0.4$ map as example, several typical query cases are shown in Fig. 11, which indicate the challenge of re-localization, with (a) lighting condition variation, (b) seasonal vegetation change, (c) dynamic objects, and (d) high-ambiguity environmental patterns. In such scenarios, the traditional handcrafted place recognition method (ORB + DBoW2) becomes almost useless, as the low-level intensities of the environment changes significantly for the cross-temporal sequences. For learning-based methods, the classic NetVLAD achieves 59.53% 10 m recall, while more modern methods (e.g., CosPlace, EigenPlaces and MixVPR) can generally achieve 85% 10 m recall. Although 85% recall seems good enough for VPR, it can’t serve as a reliable information source for the localization task, especially for unmanned systems. By naively combining the VPR results of multiple frames, the 10 m/20 m recall can increase by around 1%. In contrast, with the proposed SAS-based multi-frame VPR technique, the 10 m/20 m recall increases by approximately 10%/5%. For the SAS10 scheme, the 20 m recall achieves 100% for both ResNet50 and ResNet18 backbones. Compared to single-frame VPR methods, the introduction of SAS overcomes the challenges depicted in Fig. 11 and greatly improves the reliability of the coarse localization, which is crucial for further fine localization. Essentially, the association of spatiotemporal information serves as a cross-verification mechanism, enhancing the confidence of visual cues and leading to improved place recognition performance under significant appearance changes.
We also list the results based on a sparser map with co-visibility threshold $\Xi=0.3$. The recall is noticeably lower than the $\Xi=0.4$ case, while the EigenPlaces50 (SAS10) scheme can still achieve 92.45%/100% 10 m/20 m recall.
As extra reference, we employ the SAS technique to VPR models other than EigenPlaces. The results listed in TABLE IV verify its broad applicability and superiority over simple clustering-based multi-frame VPR.
Figure 12: Horizontal position errors of the fine localization. For SF-Loc, the $\Xi=0.4$ map is used and different multi-frame settings (1, 2, 5, 10) are employed. See TABLE IV for detailed description.
@@ VI-B2 Fine Localization
As to the evaluation of fine localization, we test the re-localization modes of ORB-SLAM3 and HLoc as the baselines, and also test GNSS-based schemes for reference, including single point positioning (SPP) and VINS-Fusion (stereo VIO + SPP).
For ORB-SLAM3 and HLoc, the coarse-to-fine localization is performed based on their own pipeline and map data (corresponding to TABLE I). For HLoc, EigenPlaces (50) is used for coarse localization and the SuperPoint + LightGlue scheme is used for fine localization.
For the proposed SF-Loc, the map data generated with different co-visibility thresholds ($\Xi=0.3/0.4$) are considered, and the coarse localization results of EigenPlaces50 (SAS10) are used. For fine localization, we test both single-frame and multi-frame modes. For multi-frame fine localization, the matching results of $N$ sequential queries are used, and the user-side relative poses are provided by stereo VIO (about 1% relative translation error, consistent with typical performance[14]). Basically, the feature matching is performed using SuperPoint + LightGlue, and the factor graph optimization implementation in Sect. V-B is used. For comparison, we also test PnP with RANSAC upon the SF-Loc map.
TABLE IV: Accuracy of the Fine Re-Localization. Horizontal Position Errors are Evaluated. $d_{\text{VPR}}$ Denotes the Median Recall Distance of Coarse Localization.
Map
$d_{\text{VPR}}$ (m)]
Method1
Avail.
@0.5 m
Avail.
@1.0 m
Avail.
@5.0 m
RMSE2
@@
- VINS-Fusion 0.00% 0.00% 42.92% 5.628
ORB [3.12]
ORB-SLAM3
0.44% 0.78% 1.00% 1.665
SP [2.42]
HLoc
89.56% 93.23% 96.34% 0.505
SF ($\Xi$=0.4)
[3.19]
PnP 72.62% 86.53% 98.15% 0.839
FGO (1) 81.53% 90.11% 98.37% 0.722
FGO (2) 90.98% 96.20% 99.35% 0.452
FGO (5) 96.09% 99.40% 99.84% 0.259
FGO (10) 98.53% 100.00% 100.00% 0.209
FGO (10)* 96.36% 99.89% 100.00% 0.219
SF ($\Xi$=0.3)
[4.30]
PnP 65.73% 83.70% 97.45% 0.934
FGO (1) 77.35% 87.29% 97.66% 0.830
FGO (2) 86.31% 93.86% 98.91% 0.614
FGO (5) 94.84% 98.26% 100.00% 0.306
FGO (10) 98.97% 100.00% 100.00% 0.198
FGO (10)* 96.54% 99.84% 100.00% 0.227
@@
    2 For RMSE calculation, only recall distance of coarse localization <<20 m and final error <<5 m cases are taken into account.
TABLE V: Accuracy of Single-Frame Fine Localization using Different Matching Methods. Using “SF-Loc, FGO(1)” Scheme ($\Xi=0.4$).
Method1
Avail.
@0.5 m
Avail.
@1.0 m
Avail.
@5.0 m
RMSE1
Time Cost
SP + LG 81.53% 90.11% 98.37% 0.722 30 + 23
AspanFormer 77.78% 87.83% 97.66% 0.882 125
RoMa 87.02% 93.59% 98.90% 0.719 371
LoFTR 79.58% 90.49% 97.50% 0.773 68
MASt3R 89.90% 94.13% 98.21% 0.580 307
MASt3R 32.54% 52.36% 95.33% 1.601 307
  • 1 For RMSE calculation, only << 5 m cases are taken into account.
    Using predicted depth instead of the depth in the map frame.
Figure 13: Examples of fine re-localization based on single-/multi-frame observations, using “SF-Loc, FGO(1)” and “SF-Loc, FGO(2)” schemes respectively. Position and attitude errors are provided in the attached boxes.
@@
As can be seen from Fig. 12 and TABLE IV, ORB-SLAM3 can hardly perform successful re-localization, as the feature matching based on local intensity pattern is greatly challenged by the significant appearance variation. For HLoc and SF-Loc, decimeter-level localization is achievable when the map matching is well performed, while outliers with several-meter error are also present, which indicates ambiguous data association. It is noted that HLoc performs better than single-frame SF-loc, which is reasonable as the HLoc map contains many more frames (see TABLE I) and stores the fine descriptors extracted from lossless raw images. In contrast, SF-Loc only uses highly compressed map frame information.
@@
It is impressive that, the localization accuracy of SF-Loc can be greatly improved by integrating multi-frame information. The examples that multi-frame observations lead to better data association and localization results are depicted in Fig. 13. As can be seen, the dynamic objects and unstable features in Fig. 13 (a), as well as the mismatches caused by repetitive textures in Fig. 13 (b), lead to significant errors in single-frame localization. The introduction of multi-frame information helps adjust the inlier set of correspondences and improve the geometric configuration, leading to more robust pose estimation. For the $\Xi=0.4$ map, SF-Loc with 10-frame FGO achieves 98.53% availability@0.5 m and an RMSE of 0.209 m. As to the absolute positioning accuracy considering the error of map frames, the availability@0.5 m is 96.36% and the RMSE is 0.219 m, which mainly satisfy the demand of decimeter-level localization accuracy. When the map becomes sparser ($\Xi=0.3$), the multi-frame localization performance remains almost unchanged, which shows the possibility of harder map compression in the cost of some redundancy.
In addition, based on the feature-free map with dense depth, detector-free matching approaches can be naturally applied in SF-Loc. In TABLE V, we compare state-of-the-art detector-free matching methods[48, 49, 50, 51] with the default SuperPoint + LightGlue scheme. The results demonstrate the possibility of achieving higher localization accuracy in cost of extra computational resources. We also consider a “map-free” localization strategy based on MASt3R’s metric depth prediction; However, its performance is relatively suboptimal, possibly due to domain discrepancies across datasets. This further underscores the value of incorporating metric-scale dense depth information into the visual structure frame map.
@@
Finally, we show the time consumption statistics of user-side localization in Fig. 14, which is tested on a laptop with Nvidia RTX 4080 Mobile GPU. With an average processing time of around 110 ms per incoming image and an input frequency of 1 Hz, the system demonstrates good real-time performance.
Figure 14: Time costs of different modules in user-side localization of SF-Loc. It is noted that VPR descriptor computation, feature extraction and matching are performed every time a frame comes and stored, so it won’t lead to significant latency during multi-frame processing. 1 Using “EigenPlaces50 (SAS10)” scheme in TABLE II. 2 Using “SF-Loc, FGO(10)” scheme in TABLE IV.
@@ VII Conclusion
In this paper, we propose a vision-centered mapping and localization framework based on the map representation of visual structure frames. The highlight of the system is the lightweight map storage by making the frame data compact and keeping the sparsity, while still being able to support high-recall, high-accuracy localization through the utilization of user-side multi-frame information.
In our future work, we will extend the system to indoor scenarios, explore hybrid map representations, and further investigate the reliability issue of visual localization.